An electrostatic force of 20N is exerted on a charge of 8.0 x 10 3 C at point P in




an electric field. What is the magnitude of the electric field intensity at point P?

1
Expert's answer
2022-02-28T11:09:39-0500

The intensity E of the electric field is "E = \dfrac{F}{q}" , so it's a force per unit charge.

"E = \dfrac{20\,\mathrm{N}}{8.0\cdot10^3\,\mathrm{C}} = 0.0025\,\mathrm{N/C \;\; (or \;\;V/m)}."